Fermi National Laboratory
 

Operation Summary for April 6, 2011 - April 8, 2011

 

Wednesday April 6

The day shift began with the Tevatron (TEV) off and in access to investigate a sump pump and vacuum pump problem, with Operations monitoring the Recycler stash (442.0E10), with the Antiproton Source (Pbar) stacking (12.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, the TEV A-E sectors and F-Sector accesses ended at 12:47 PM. Operators began the TEV turn-on. Proton Source RF issues lasted all shift.

During the evening shift, Operations put the TEV into HEP shot setup and established store 8629 at 6:53 PM with an average initial luminosity (AIL) of 313.8E30. Operators accessed the MI-12A enclosure to investigate a closed beam valve at 7 PM.

Thursday April 7

The midnight shift began with Operations monitoring store 8629 (141.29E30) and the stash (171.1E10), with Pbar stacking (32.8E10), and with MiniBooNE taking beam.

During the midnight shift, operators reported that the MI ramp tripped off at 6:45 AM on a voltage-to-ground (ground fault) indication at MI-60N; this trip caused the loss of one transfer of antiprotons to the Recycler.

During the day shift, Operations terminated store 8629 at 8:39 AM, put the TEV into HEP shot setup, and established store 8630 at 9:50 AM with an AIL of 328.59E30. Operations established beam to the returning Meson FTFB T-978 (CALICE) experiment at 3:46 PM.

During the evening shift, Operations sent low intensity and low repetition rate beam to NuMI for target scans at 4:56 PM. Operations disabled beam to T-978 at 6:04 PM.

Friday April 8

The midnight shift began with Operations monitoring store 8630 (78.3E30) and the stash (352.2E10), with Pbar stacking (17.8E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8630 at 1:41 AM, put the TEV into shot setup, and established store 8631 at 3 AM with an AIL of 344.09E30. Operations established beam to T-978 around 4:15 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct studies and also allow more NuMI target scans. NuMI expects to resume normal operations on Saturday, 4/9/11.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/08/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory